# An Indoor Radio Mapping Dataset Combining 3D Point Clouds and RSSI

### v2

@@ Abstract
The growing number of smart devices supporting bandwidth-intensive and latency-sensitive applications, such as real-time video analytics, smart sensing, Extended Reality (XR), etc., necessitates reliable wireless connectivity in indoor environments. In such environments, accurate design of Radio Environment Maps (REMs) enables adaptive wireless network planning and optimization of Access Point (AP) placement. However, generating realistic REMs remains difficult due to the variability of indoor environments and the limitations of existing modeling approaches, which often rely on simplified layouts or fully synthetic data. These challenges are further amplified by the adoption of next-generation Wi-Fi standards, which operate at higher frequencies and suffer from limited range and wall penetration. To support the efforts in addressing these challenges, we collected a dataset that combines high-resolution 3D LiDAR scans with Wi-Fi RSSI measurements collected across 20 setups in a multi-room indoor environment. The dataset includes two measurement scenarios, the first without human presence in the environment, and the second with human presence, enabling the development and validation of REM estimation models that incorporate physical geometry and environmental dynamics. The described dataset supports research in data-driven wireless modeling and the development of high-capacity indoor communication networks.
keywords:
RSSI, Point cloud, REM, Indoor Radio, Wi-Fi
@@ Methods
ine LiDAR X-coordinate (m) Y-coordinate (m) Z-coordinate (m)
ine Avia1 1.38 7.34 0.95
ine Avia2 1.70 10.33 0.97
ine Avia3 4.85 10.93 1.05
ine Avia4 8.24 8.73 1.08
ine Avia5 7.73 0.14 1.57
ine Avia6 3.81 0.24 1.44
ine
@@
ine LiDAR X-coordinate (m) Y-coordinate (m) Z-coordinate (m)
ine Avia1 0.00 1.00 1.30
ine Avia2 3.70 16.95 1.30
ine Avia3 32.20 1.00 1.30
ine VLP1 3.70 0.50 1.30
ine VLP2 2.00 8.75 1.30
ine
@@
(a) Point cloud of the indoor environment considered for the measurements, orange=office, blue=corridor.
(b) RSSI measurements points and contour plot based on the measurements of setup 6. Blue=lower, Yellow=higher
@@
The RSSI data is available in two separate formats. The first one is a .csv file (visualized in Table 3) containing the raw measurements of the RSSI, together with other AP-specifics, such as estimated distance to the AP, exact frequency, channel, etc. In another .h5 file, contents visualized in Section HDF5 File Structure: WiFi_RSSI_data2025-05-21_15-51-16.h5, where the extracted RSSI measurements (data) are in matrix format as visualized in Figures 6b and 7, allowing for direct operation on the RSSI data. Furthermore, labels, regarding the setup number (setup), AP location (ap-locations), and measurement flow (indices) are also included in the .h5 file. Regarding the potential reuse, the code for extracting the RSSI data and some simple manipulations, such as visualization of particular setups and filling missing values, will also be provided as part of the toolbox. Missing data appear mostly on the outermost data acquisition points, such as 50-53 in Figure 6b, due to the excessive distance and multiple walls between the AP and UE.
@@
Another notable phenomenon is the observation of relatively weaker or comparable RSSI values at the AP’s exact location, compared to immediately adjacent positions. For example, in setup 2 (Figure 7c), the AP is placed at location 3, where an RSSI of 33dBm-33dBm is recorded. Interestingly, the neighboring location 2 shows a significantly stronger signal of 19dBm-19dBm. This effect may be attributed to the vertical proximity of the measurement device to the AP, capturing a point within the antenna’s radiation null. Omnidirectional antennas typically exhibit a "donut-shaped" radiation pattern, which results in weaker signal strength directly above or below the AP.
@@
(a) People not present.
(b) 7-10 people present.
(c) Weaker RSSI above the AP
(d) Stronger RSSI in neighboring locations.

### v3

@@ Abstract
The growing number of smart devices supporting bandwidth-intensive and latency-sensitive applications, such as real-time video analytics, smart sensing, Extended Reality (XR), etc., necessitates reliable indoor wireless connectivity. In such environments, accurate Radio Environment Maps (REMs) enable adaptive wireless network planning and optimization of Access Point (AP) placement. However, generating realistic REMs remains difficult due to the variability of indoor environments and the limitations of existing modeling approaches, which often rely on simplified layouts or synthetic data. These challenges are further amplified by the adoption of next-generation Wi-Fi standards, operating at higher frequencies with limited range and wall penetration. To support progress in this area, we collected a dataset that combines high-resolution 3D LiDAR scans with Wi-Fi RSSI measurements across 20 setups in a multi-room indoor environment. It includes two measurement scenarios, one with and one without human presence, enabling development and validation of REM estimation models that incorporate physical geometry and environmental dynamics. The described dataset supports research in data-driven wireless modeling and the development of high-capacity indoor communication networks.
Keywords: indoor radio mapping, LiDAR, RSSI, wireless, dataset
@@ Methods
LiDAR X-coordinate (m) Y-coordinate (m) Z-coordinate (m)
Avia1 1.38 7.34 0.95
Avia2 1.70 10.33 0.97
Avia3 4.85 10.93 1.05
Avia4 8.24 8.73 1.08
Avia5 7.73 0.14 1.57
Avia6 3.81 0.24 1.44
@@
LiDAR X-coordinate (m) Y-coordinate (m) Z-coordinate (m)
Avia1 0.00 1.00 1.30
Avia2 3.70 16.95 1.30
Avia3 32.20 1.00 1.30
VLP1 3.70 0.50 1.30
VLP2 2.00 8.75 1.30
@@
The RSSI data is available in two separate formats. The first one is a .csv file (visualized in Table 3) containing the raw measurements of the RSSI, together with other AP-specifics, such as estimated distance to the AP, exact frequency, channel, etc. In another .h5 file, with contents visualized at the end of this section, the extracted RSSI measurements (data) are in matrix format as visualized in Figures 6b and 7, allowing for direct operation on the RSSI data. Furthermore, labels, regarding the setup number (setup), AP location (ap-locations), and measurement flow (indices) are also included in the .h5 file. Regarding the potential reuse, the code for extracting the RSSI data and some simple manipulations, such as visualization of particular setups and filling missing values, will also be provided as part of the toolbox. Missing data appear mostly on the outermost data acquisition points, such as 50-53 in Figure 6b, due to the excessive distance and multiple walls between the AP and UE.
@@
Another notable phenomenon is the observation of relatively weaker or comparable RSSI values at the AP’s exact location, compared to immediately adjacent positions. For example, in setup 2 (Figure 7c), the AP is placed at location 3, where an RSSI of 33dBm-33dBm is recorded. Interestingly, the neighboring location 2 shows a significantly stronger signal of 19dBm-19dBm. This effect may be attributed to the vertical proximity of the measurement device to the AP, capturing a point within the antenna’s radiation null. Omnidirectional antennas typically exhibit a ”donut-shaped” radiation pattern, which results in weaker signal strength directly above or below the AP.
@@
Data availability
The data produced during this study is publicly available on Zenodo https://doi.org/10.5281/zenodo.15791300, in the data folder. It consists of the 3D point clouds in three separate ply files, which include the measurement area, consisting of two point clouds, and one example output of the point cloud registration, produced by the example_operations.ipynb script. The RSSI data is provided in a concise csv file and in raw h5 format. The repository organization is shown in Figure 8.
Figure 8: Organization of the dataset repository.